OpenStudy (anonymous):

PLEASE HELP I WILL FAN AND MEDAL!!!! Find a number that is between 8/11 and 0.85, 85 repeating. A.157/198 B.88/99 C.23/33 D.39/66

ganeshie8 (ganeshie8):

Notice \(\dfrac{8}{11} = \color{blue}{\dfrac{72}{99}}\) and \(0.\overline{85} = \dfrac{85}{10^2-1} = \color{blue}{\dfrac{85}{99}}\)

OpenStudy (anonymous):

right I figured that out for 0.85 but how did \[\frac{ 8 }{ 11 }\] become \[\frac{ 72 }{ 99 }\]

ganeshie8 (ganeshie8):

Multiply top and bottom of \(\frac{8}{11}\) by \(9\)

OpenStudy (anonymous):

Oh okay I understand better now so what do i do next?

ganeshie8 (ganeshie8):

Even better, change the denominator of everything to \(198\) so that comparison becomes easy (99*2 = 198)

ganeshie8 (ganeshie8):

A.157/198 = `157/198` B.88/99 = `176/198` C.23/33 = `138/198` D.39/66 = `117/198`

ganeshie8 (ganeshie8):

You want to find a number between \(\dfrac{8}{11} = \dfrac{72}{99}=\color{blue}{\dfrac{144}{198}}\) and \(0.\overline{85} = \dfrac{85}{10^2-1} = \dfrac{85}{99}= \color{blue}{\dfrac{170}{198}}\)

OpenStudy (anonymous):

It would be \[\frac{ 157 }{ 198 }\] I think @ganeshie8

ganeshie8 (ganeshie8):

Yep!

OpenStudy (anonymous):

oh my gosh thank you so much

ganeshie8 (ganeshie8):

yw:)